There are 480 welders employed at the Weller Shipyards Corporation. A sample of 45 welders revealed that 27 had graduated from a registered welding course. Construct the 99% confidence interval for the proportion of all welders who graduated from a registered welding course. (Round the final answers to 3 decimal places.)
Answer #1

Solution :

Given that,

Point estimate = sample proportion = \hat p = x / n = 27 / 45 = 0.600

1 - \hat p = 1 - 0.600 = 0.4

Z\alpha/2 = 2.576

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 2.576 * (\sqrt((0.600 * 0.4) / 45)

Margin of error = E = 0.188

A 99% confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.600 - 0.188 < p < 0.600 + 0.188

0.412 < p < 0.788

The 99% confidence interval for the population proportion p is : 0.412 , 0.788